Скачать презентацию Virgo commissioning update VSR 1 first month summary Скачать презентацию Virgo commissioning update VSR 1 first month summary

e891186e0fb8fe61ed40f4a178643227.ppt

  • Количество слайдов: 28

Virgo commissioning update VSR 1 first month summary E. Tournefier WG 1 meeting, Cascina Virgo commissioning update VSR 1 first month summary E. Tournefier WG 1 meeting, Cascina June 20 th , 2007

Activities since last WG 1 meeting (March 28 th) • Virgo complete electrical shutdown Activities since last WG 1 meeting (March 28 th) • Virgo complete electrical shutdown (April 6 th -> 18 th) Infrastructure works: – Electrical works: UPS replacement, reorganisation, … – Computing: civil work, re-organisation, new file server, … Detector improvements - Acoustic enclosure around external detection bench - Replacement of Brewster window with a larger one Restart went ok: ITF relocked within ~2 days • 1 month of commissioning before the run – Acoustic enclosures at end benches – Diffused light investigations + some noise reduction – Improvements of controls (longitudinal, angular, suspensions) – Change of naming convention (V 1: channel_name)for easier data exchange with LSC • VSR 1 started on May 18 th as planned – Short commissioning breaks for improvements / problem fixing

Acoustic enclosures • External detection bench: – Displace bench from the tower (~50 cm) Acoustic enclosures • External detection bench: – Displace bench from the tower (~50 cm) – Install acoustic panels (a room into the room) • End benches: – Simpler: build a room around the benches End buildings Detection

Acoustic noise reduction: end benches - Previous actions: Improvement of the optical setup But: Acoustic noise reduction: end benches - Previous actions: Improvement of the optical setup But: projection of acoustic noise of end benches was still close to the measured sensitivity Tentative acoustic noise projection before Þ Acoustic enclosure needed Result: impact reduced above 100 Hz Tentative acoustic noise projection with acoustic enclosu Seismic noise on the optical bench - No acoustic enclosure - With acoustic enclosure

New Brewster window Why change the detection Brewster window: - Brewster-detection tower was shown New Brewster window Why change the detection Brewster window: - Brewster-detection tower was shown to be an area very sensitive to acoustic noise - tentative noise projections: Brewster could be a limiting factor - old Brewster didn’t respect the rule: diameter > 5 x beam waist Tentative projection of ‘Brewster noise’ Detection tower Þ Replace Brewster with a larger one (April shutdown) Followed by diffused light investigations No obvious improvement in sensitivity See Raffaele’s talk Brewster link

Other actions for noise reductions (last WG 1 meeting) BS actuator noise reduced Replaced Other actions for noise reductions (last WG 1 meeting) BS actuator noise reduced Replaced pump of SR tower with a smaller one (noise reduced but path is unclear)

Mirror actuator noise reduction • Electronic noise of the coil drivers + DAC Þ Mirror actuator noise reduction • Electronic noise of the coil drivers + DAC Þ Shaping filters, more gain to reduce its impact z. Corr I (A) to mirror coil ÞBS actuator noise reduced in May ÞThe arm mirrors actuators need to be improved can be done during VSR 1

Actuators calibration Simple cavity filtering (500 Hz pole) Actuation TFs Simple pendulum More precise Actuators calibration Simple cavity filtering (500 Hz pole) Actuation TFs Simple pendulum More precise measurement of the actuators response Þ frequency dependence could be due to Eddy current effects Þ Change of 20% at high frequencies 10%

See Bastian’s talk Longitudinal controls Usual noise subtraction (online): -MICH and PRCL corrections sent See Bastian’s talk Longitudinal controls Usual noise subtraction (online): -MICH and PRCL corrections sent to the end mirrors to subtract the noise reintroduced by these loops into the dark fringe - coefficient is frequency dependent MICH noise reduction ~ 50 B 2_ACpmix PRCL MICH - with =0 - with measured (f)

Longitudinal controls: noise budget - B 1_ACp - BS control noise - PR control Longitudinal controls: noise budget - B 1_ACp - BS control noise - PR control noise BS control noise limits the sensitivity below 40 Hz => What is the origin of this noise?

Origin of BS control noise ? • Related with acoustic noise inside the laser Origin of BS control noise ? • Related with acoustic noise inside the laser lab Þ Beam jitter seen by B 2 photodiode (used for PR & BS)? • Sometimes disappears … but acoustic noise does not change Input beam jitter Beam jitter vs dark fringe Acoustic noise in laser lab Beam jitter vs acoustic noise

Angular controls • Second modulation frequency (8 MHz) now used for Q 22 => Angular controls • Second modulation frequency (8 MHz) now used for Q 22 => Get rid of DC signal • Control loops improvements • Mirror centering ~ 1 mm Angular noise budget -Angular controls do not limit the sensitivity - Reduction of electronic/shot noise in error signal still possible

Suspension control: z -> couplings Compensation of non-linear coupling: Large seismic activity Þ large Suspension control: z -> couplings Compensation of non-linear coupling: Large seismic activity Þ large correction sent to marionetta (z. M) Þ y motion of F 7 and payload Þ Side effects on alignment Solution: compensate for z -> y coupling Need a quadratic compensation: y=c*z. M +d*z. M 2 Dark fringe -No quadratic compensation -With quadratic compensation

Suspension control: top stage control Old problem: environmental noise ( seism) reintroduced by Lvdt Suspension control: top stage control Old problem: environmental noise ( seism) reintroduced by Lvdt used for top stage control => Use the ITF longitudinal error signal instead (Global Inverted Pendulum Control) => z correction to the mirror well reduced GC (reconstructed z) VSR 1: enabled only for NE-WE No GIPC With GIPC LVDT

Sensitivity vs wind and sea activity Matteo’s empirical formula: Horizon = H 0(1 - Sensitivity vs wind and sea activity Matteo’s empirical formula: Horizon = H 0(1 - a x seism – b x wind) WSR 1 (Sept 2006) VSR 1 (May 2007) a= 0. 37, b= 0. 37 => Sensitivity to bad weather conditions is well reduced a= 0. 07, b= 0. 04

Lock acquisition: thermal transient troubles • Variations in thermal transient : seems related to Lock acquisition: thermal transient troubles • Variations in thermal transient : seems related to activities in laser lab Sometimes need to wait few hours before being able to lock

Thermal effects: sidebands image • Image of the sidebands (first step of phase camera): Thermal effects: sidebands image • Image of the sidebands (first step of phase camera): scanning system on B 5 beam: - the beam is scanned on a pin-hole + photodiode - demodulated signal is used to reconstruct the sidebands shape Waiting for thermal compensation system + complete phase camera

VSR 1 noise budget Still magnetic noise? Mystery noise VSR 1 noise budget Still magnetic noise? Mystery noise

Mystery noise • Some observations: – Coupling to DF increases with misalignments – Sudden Mystery noise • Some observations: – Coupling to DF increases with misalignments – Sudden decreases/increases (also seen in frequency noise) – Structures not excited with noise injections/tapping tests around benches/towers Keep searching…

VSR 1: first month summary Thanks to Roberto for the material VSR 1: first month summary Thanks to Roberto for the material

Run organisation • Goal: – Take science data for a minimum period of 4 Run organisation • Goal: – Take science data for a minimum period of 4 months starting from the 18 th May 2007 at 23: 00 (LT) • Shifts have been planned until 20 th September 2007: • 18 weeks, 378 shifts. – Shift team and time schedule: • Team: 1 operator + 1 physicist, • Shift period: 8 hours • Shift start times: 07: 00 – 15: 00 – 23: 00 Permanent ON Call: – 11 Virgo ITF Sub-systems: • (Alignment, DAQ, Detection, Electronics, Locking, Global Control, Laser &Injection, MSC, On line Computing &Network, Software, Suspension Electronics & Software) – 4 EGO Site Services: (Network & Global Security System, Site Infrastructure-Conditioning, Site Infrastructure-Electricity, Vacuum)

Duty cycle – Duty cycle in science mode ~ 85% – 42 unlocks from Duty cycle – Duty cycle in science mode ~ 85% – 42 unlocks from “Science Mode” registered up to 17/06/07 – Longest Lock: about 58. 5 hours , Longest Unlock: about 16. 1 hours

Science mod recovery time – Longest unlock: 16 hours – Smallest ~ 30 mn Science mod recovery time – Longest unlock: 16 hours – Smallest ~ 30 mn (time needed for lock acquisition)

Lock duration – Longest lock: 58 hours Lock duration – Longest lock: 58 hours

Unlocks • Analysis of unlocks ongoing (preliminary statistics) Main unlock reasons: - IMC fast Unlocks • Analysis of unlocks ongoing (preliminary statistics) Main unlock reasons: - IMC fast unlocks Partially cured (loop gain reduced) - Global control crashes Pb = longer recovery (susp excited) - Earthquakes Manual unlocks for: - Weekly maintenance - Commissioning breaks Only 3 unlocks due to bad weather

Commissioning breaks Alignment drifts observed during locks: related to Q 81 quadrant mis-centering (used Commissioning breaks Alignment drifts observed during locks: related to Q 81 quadrant mis-centering (used for BS control) => spoils error signal (AC) Þ Temporary solution: realign quadrant from time to time Þ Improvement: include Q 81 asymmetry in the error signal for BS control Small drifts still there Þ Try to use another error signal (B 1 p? ) or Þ Install galvanometer to keep centering in science mode IMC Fast unlocks - Reduction of IMC loop gain 1 unlock per couple of day => ~ 1 per week - more investigations ongoing (electronics) Suspension controls: some improvements /tests Dark fringe Q 81 asymmetry realignment

Long term observations: Etalon effect! • Etalon effect: small FP cavity inside input mirrors Long term observations: Etalon effect! • Etalon effect: small FP cavity inside input mirrors (due to AR face) => effective reflectivity modulated with mirror thickness (temperature) -FP transmitted power -Input mirror temperature FP Transmitted power B 7/B 8 phd Temperature variation

Some short term plans • Continue noise investigations – ‘Mystery noise’ ( 200 Hz) Some short term plans • Continue noise investigations – ‘Mystery noise’ ( 200 Hz) – Magnetic noise (50 - 100 Hz) – Low frequency noise in laser lab (=> BS error signal) – Reduce mirror actuator noise (cavity mirrors) • Controls: – Alignment drifts: get rid of Q 81 or use galvanometer – Suspension control improvements in preparation